A 5.a7g bullet leaves the muzzle of a rifle with a speed of 334m/s. What force (assumed constant) is exerted on the bullet while it is traveling down the 0.842 m long barrel of the rifle?
Answer #1

using , v²-0²=zad (334) - 0 (2 x ax 0.842) a = 66244.655 m/ F = ma = 7 * 66244.655 1000 (463.7N Force eccited on bullet (463.please ask your doubts or queries in the comment section below.

Please kindly upvote if you are satisfied with the solution.

Thank you.